Insourcing incurs a periodic fixed cost of $20,000 and a $4.50 variable cost. Outsourcing incurs a periodic cost of $45,000 and a $2.50 variable cost. Over what ranges of demand is each option best?
Answer #1

Given that insourcing has higher level of variable cost and lower fixed cost, we can easily see that insourcing is the better choice for numbers below the crossover point. A crossover point is a value or number of units where both the options are equal. Thus solving for crossover point we get

20000 + 4.5X = 45000 + 2.5X

X = (45000-20000)/(4.5-2.5)= 12500

This means the correct answer is

Outsourcing is optimal for demand > 12500. Insourcing is optimal for demand < 12500